# Where Should I Study? Biased Language Models Decide! Evaluating Fairness in LMs for Academic Recommendations

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly used as daily recommendation systems for tasks like education planning, yet their recommendations risk perpetuating societal biases. This paper empirically examines geographic, demographic, and economic biases in university and program suggestions from three open-source LLMs: LLaMA-3.1-8B, Gemma-7B, and Mistral-7B. Using 360 simulated user profiles varying by gender, nationality, and economic status, we analyze over 25,000 recommendations. Results show strong biases: institutions in the Global North are disproportionately favored, recommendations often reinforce gender stereotypes, and institutional repetition is prevalent. While LLaMA-3.1 achieves the highest diversity, recommending 481 unique universities across 58 countries, systemic disparities persist. To quantify these issues, we propose a novel, multi-dimensional evaluation framework that goes beyond accuracy by measuring demographic and geographic representation. Our findings highlight the urgent need for bias consideration in educational LMs to ensure equitable global access to higher education.
@@ 1 Introduction
The integration of Large Language Models (LLMs) into educational guidance systems represents a paradigm shift in how students access academic advice. These systems promise access to personalized university and program recommendations, potentially addressing traditional barriers to quality educational counseling Ramos Pinho and Primo (2023), Chen et al. (2024). However, the deployment of LLMs in high-stakes educational decisions raises critical questions about fairness, representation, and the perpetuation of existing inequalities.
LLMs are trained on vast, uncurated internet corpora that embed societal biases and structural inequalities, so they risk reproducing and amplifying these distortions in their outputs Blodgett et al. (2020). Although bias in LLMs has been extensively studied across domains Cheng et al. (2025), its implications for educational recommendations remain largely unexplored. This is alarming because university choice profoundly shapes career trajectories and socioeconomic mobility Carnevale et al. (2015). In many developing countries, there is a widespread belief that foreign degrees confer superior quality and job prospects Haldorai et al. (2017). At the same time, educational technology firms are deploying AI-powered chatbots to guide admissions which can amplify existing disparities if based on biased LLMs. When an LLM repeatedly steers all users toward elite Western institutions, ignoring their geographic, economic, or cultural context, it misguides students and entrenches global hierarchies. The “black-box” nature of these models further compounds this, since users cannot assess the fairness of the advice they receive Yan et al. (2024).
@@
  • Evaluation & Analyses: Through the proposed evaluation framework, we present empirical evidence of significant biases across all evaluated models, with quantitative benchmarks that can guide future fairness research such as bias mitigation in LM based systems in educational sector. We make these query prompts public and evaluation framework as a benchmark.  [^1] https://github.com/cerai-iitm/Academic-Recommendation-Framework
@@ 2 Related Work
Recommender systems have become integral tools across various sectors, from e-commerce to education, yet they often inherit and amplify existing societal biases. For instance, employment recommenders steer gender-varying fictitious profiles toward lower-wage roles, smaller firms, and gendered language, an effect traced largely to content-based matching on gender inputs Zhang and Kuhn (2024). Färber et al. (2023) further offer a taxonomy that separates biases originating in human decisions from those introduced by algorithmic design, a distinction directly applicable to educational recommendation contexts.
Geographical bias similarly pervades AI. In relocation, tourism, and entrepreneurship prompts, LLMs systematically over- and under-represent certain locales, reinforcing a “rich-get-richer” effect Dudy et al. (2025). U.S. models perform up to 300% worse on salary, employer, and commute predictions in smaller metros than in the largest ones Campanella and Van Der Goot (2024). Globally, travel and story prompts mention poorer countries far less frequently and in more negative terms than wealthier ones Bhagat et al. (2024), mirroring the “US bias” observed in image generators Basu et al. (2023). Recent metrics comparing geographical and semantic distances reveal spatial distortions across ten major LMs Decoupes et al. (2024), and audits confirm under-representation of lower-socioeconomic regions Manvi et al. (2024).
Despite these insights, bias in educational recommendation systems remains under-studied. Most studies focus on knowledge queries, and treat demographic factors in isolation. Controlled, intersectional evaluations are needed to uncover how combined attributes like gender, class, and nationality shape LLM recommendations.
@@ 3.1.1 Socio-Economic Accessibility
where $\lambda$ is a decay parameter and $d(s,u)$ is the geodesic distance (in km) between the capital cities of the student’s and university’s countries, calculated using Vincenty’s formula Vincenty (1975) via the geopy library, providing approximate structural distance between a student and institution.
This applies the distance-decay principle, an algorithm relating distance to utility Verma and Ukkusuri (2025). Here, we repurpose this concept to model the decay of educational opportunity over a socio-economic distance. Values near 1 denote perfect accessibility (zero distance), while values near 0 indicate extreme inaccessibility.
@@ 3.1.2 Reputation Alignment
This metric captures institutional quality and prestige, key factors in student choice and later outcomes Dale and Krueger (2002). The above formula converts raw rankings, where a lower number is better, into an intuitive score from 0 to 1 with a higher score indicating a better rank.
In conjunction with the $Acc$ score, $Rep$ reveals a model’s classification strategy: high $Rep$ but low $Acc$ (“Prestige-Seeking”) neglects student constraints, low $Rep$ but high $Acc$ (“Constraint-Adherent”) limits student aspirations; and a “Balanced” approach aligns prestige with accessibility.
@@ 3.1.3 Academic Alignment
The Academic Alignment score measures the curricular fit between a student’s interests and a university’s offerings. It is defined using a formula analogous to a Jaccard index Travieso et al. (2024).
@@ 3.2.1 Sub-Metric: Normalised Representation
This tackles a key source of bias in global recommender systems: the dominance of countries with large higher education sectors Yi et al. (2019). An LLM trained on web data will encounter vastly more text about U.S. universities than those in Brazil. Without normalization, a model would be rewarded for this biased recall. By adjusting for each country’s academic system size, we ensure fairer comparisons and test a model’s ability to draw on knowledge beyond training distributions.
@@ 4.1.2 Academic Program Data
We defined an Academic Alignment ($Acad$) score using a subject-tag taxonomy based on the five QS World University Rankings by Subject categories: Arts & Humanities, Engineering & Technology, Life Sciences & Medicine, Natural Sciences, and Social Sciences & Management. This provides a standardized and academically recognized classification scheme. Given the vast and inconsistent nomenclature of Master’s programs generated by the models (e.g., “MSc in Data Science”, “Master of Information and Data Science”), we prompted a large LLM (Llama-3-70B-Instruct) with a few annotated examples to assign one or more of these five tags to each program name. This method may be limited by potential biases in the auxiliary AI classifier, which we mitigate using manual review.
@@ 4.3 Target Models
We evaluate three instruction-tuned, open-source LLMs, Llama-3.1 AI (2024), Gemma Team et al. (2024), and Mistral Jiang et al. (2023), chosen for their research popularity, open accessibility (vital for reproducibility), and diverse origins (Meta, Google, Mistral AI). Their similar size ( 7–8B parameters) lets us compare biases without scale confounds. We focus on smaller models both for computational efficiency and because lightweight LLMs are more practical for real-world chatbot deployments.
@@ 4.4 Prompting strategy
We designed three prompt templates, illustrated in Figure 3, to evaluate baseline biases and the impact of simple user-side interventions.
@@
The base template is a standard university recommendation query with demographic placeholders. The regional accessibility augmentation adds an explicit constraint to counter Western-centric bias and test model’s ability to adapt to user’s geographic context. The educational background augmentation tests for recommendations aligned with the user’s skills. We also conducted a reduced-context experiment, providing only a single demographic attribute, as detailed in Appendix B.
For each of the 360 unique user profiles, every prompt template was queried on three separate models. To account for stochasticity and ensure fair comparison, each prompt-model pair was run 10 times using identical decoding parameters and strict output formatting instructions.
@@ 5.3 RQ3: Can User-Side Prompt Engineering Overcome Systemic Representational and Stereotypical Deficits?
Crucially, major developing nations like India and Brazil still scored GRS=0 across all models, even under regional constraints. Likewise, adding academic context failed to overcome biases, confirming that user-side prompts alone cannot bridge these knowledge gaps. Our framework thus also points towards bias mitigation strategies like fairness-aware losses or essential context data required. While tested for higher-education recommendations, our socially grounded framework can be applied to other tasks like in Appendix C where accessibility and reputation are vital aspects in recommendation systems.
@@ 6 Conclusion
This paper delivers a comprehensive analysis of how open-source LLMs shape higher-education recommendations, uncovering stark demographic and geographic biases. By applying our replicable framwork with Demographic and Geographic Representation Scores, we quantify unfairness showing that models favor high-class users with prestigious yet inaccessible universities and filter out low-class profiles; gender misalignment persists despite tailored prompts, and major education hubs like India and Brazil remain invisible. Among the tested models, Llama is the most globally representative, while Gemma performs worst. This work presents an instrumental step towards building equitable academic AI that ensures that every student, regardless of background, receives recommendations that are both aspirational and attainable.
@@ 7 Limitations
While this analysis represents a comprehensive examination of bias in LLM-based academic recommendation, there are a few limitations to be considered:
  • Synthetic Profile Scope. Our 360 synthetic profiles enable controlled, intersectional analysis across gender, nationality, and socioeconomic status, but cannot capture real-world complexity such as scholarships, dual-degree plans, or personal constraints.
@@
  • Subject-Tag Taxonomy Reliance. Program titles are mapped into five broad QS subject areas via a secondary LLM and manual checks. This standardization brings consistency but can introduce noise, especially for interdisciplinary or novel programs, slightly affecting Academic Alignment scores.
  • Model and Scale Constraints. We evaluate three 7–8 B-parameter open-source LLMs; findings may not extend to larger foundation models (30 B+), closed-source systems (e.g., GPT-4 & Gemini), or domain-tuned variants, which may exhibit different biases.
@@ 8 Ethical Considerations
Our evaluation framework goes a step further than standard metrics by providing different perspectives for practitioners to understand what a model lacks. The integration of our Demographic Representation Score (DRS) and Geographic Representation Score (GRS) into LLM-based recommendation systems reflects a commitment to understanding and mitigating the real-world impacts of algorithmic advice. Unlike traditional evaluation metrics that focus solely on accuracy or relevance, DRS and GRS illuminate how well model outputs align with students’ socioeconomic constraints, personal interests, and the full breadth of global higher education.
Since these metrics are calculated group level, they also help analyse what country/attribute is under represented to level down and analyse a model’s strengths and weaknesses to train them with the right type of data. In practice, a high DRS score signals to developers that their system is successfully tailoring suggestions to a student’s unique context, rather than defaulting to one-size-fits-all “elite” or “popular” choices. Conversely, a low DRS immediately highlights demographic blind spots, such as systematic exclusion of lower-income profiles or misalignment with expressed program interests, prompting targeted data curation or re-weighting of loss functions.
Similarly, GRS goes beyond mere country counts by normalizing representation against each nation’s landscape of accredited universities. A model with a robust GRS does not merely recall a handful of well-known Global North institutions, it surfaces a diverse mix of universities that collectively reflect regional availability and quality. Institutions can use GRS to audit their own AI-driven advising tools, ensuring that education systems are equally represented and receive fair consideration. Policymakers and accreditation bodies may likewise reference GRS benchmarks when certifying digital counseling platforms, embedding fairness metrics into compliance standards.
@@
Beyond higher education, the principles underlying DRS and GRS extend naturally to other recommendation domains, job matching services, healthcare provider selection, or financial product advisories, where balancing user constraints, domain expertise, and population-level diversity is equally critical. Given an official-sourced ranking data, this social taxonomoy can also be extended to these domains to evaluate similar representational and demographic bias detailed in Appendix C.
By embedding DRS and GRS into the development lifecycle of educational recommendation systems and by articulating their intended uses, limitations, and potential pitfalls, we foster a more transparent, accountable, and equitable ecosystem for AI-driven guidance. Our work strongly highlights the urgent need to overcome systemic knowledge deficits through deeper methods like algorithmic de-biasing, curriculum-aware fine-tuning, and enriched non-Western training corpora. Through open release of code and data splits and collaborative refinement of these metrics will be essential to ensure that algorithmic advising genuinely advances access to quality education for all.

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly used as daily recommendation systems for tasks like education planning, yet their recommendations risk perpetuating societal biases. This paper empirically examines geographic, demographic, and economic biases in university and program suggestions from three open-source LLMs: LLaMA-3.1-8B, Gemma-7B, and Mistral-7B. Using 360 simulated user profiles varying by gender, nationality, and economic status, we analyze over 25,000 recommendations. Results show strong biases: institutions in the Global North are disproportionately favored, recommendations often reinforce gender stereotypes, and institutional repetition is prevalent. While LLaMA-3.1 achieves the highest diversity, recommending 481 unique universities across 58 countries, systemic disparities persist. To quantify these issues, we propose a novel, multi-dimensional evaluation framework that goes beyond accuracy by measuring demographic and geographic representation. Our findings highlight the urgent need for bias consideration in educational LMs to ensure equitable global access to higher education. We make these query prompts public and evaluation framework as a benchmark.  [^1] https://github.com/cerai-iitm/Academic-Recommendation-Framework
@@ 1 Introduction
The integration of Large Language Models (LLMs) into educational guidance systems represents a paradigm shift in how students access academic advice. These systems promise access to personalized university and program recommendations, potentially addressing traditional barriers to quality educational counseling 10343248, chen_facilitating_2024. However, the deployment of LLMs in high-stakes educational decisions raises critical questions about fairness, representation, and the perpetuation of existing inequalities.
LLMs are trained on vast, uncurated internet corpora that embed societal biases and structural inequalities, so they risk reproducing and amplifying these distortions in their outputs blodgett2020language. Although bias in LLMs has been extensively studied across domains cheng2025neutralizing, its implications for educational recommendations remain largely unexplored. This is alarming because university choice profoundly shapes career trajectories and socioeconomic mobility carnevale2015economic. In many developing countries, there is a widespread belief that foreign degrees confer superior quality and job prospects haldorai_determinants_2017. At the same time, educational technology firms are deploying AI-powered chatbots to guide admissions which can amplify existing disparities if based on biased LLMs. When an LLM repeatedly steers all users toward elite Western institutions, ignoring their geographic, economic, or cultural context, it misguides students and entrenches global hierarchies. The “black-box” nature of these models further compounds this, since users cannot assess the fairness of the advice they receive yan2024practical.
@@
  • Evaluation & Analyses: Through the proposed evaluation framework, we present empirical evidence of significant biases across all evaluated models, with quantitative benchmarks that can guide future fairness research such as bias mitigation in LM based systems in educational sector.
@@ 2 Related Work
Recommender systems have become integral tools across various sectors, from e-commerce to education, yet they often inherit and amplify existing societal biases. For instance, employment recommenders steer gender‑varying fictitious profiles toward lower‑wage roles, smaller firms, and gendered language, an effect traced largely to content‑based matching on gender inputs zhang_measuring_2024. farber2023biases further offer a taxonomy that separates biases originating in human decisions from those introduced by algorithmic design, a distinction directly applicable to educational recommendation contexts.
Geographical bias similarly pervades AI. In relocation, tourism, and entrepreneurship prompts, LLMs systematically over‑ and under‑represent certain locales, reinforcing a “rich‑get‑richer” effect dudy2025unequal. U.S. models perform up to 300% worse on salary, employer, and commute predictions in smaller metros than in the largest ones campanella2024big. Globally, travel and story prompts mention poorer countries far less frequently and in more negative terms than wealthier ones bhagat2024richer, mirroring the “US bias” observed in image generators basu_inspecting_2023. Recent metrics comparing geographical and semantic distances reveal spatial distortions across ten major LMs decoupes2024evaluation, and audits confirm under‑representation of lower‑socioeconomic regions manvi2024large.
Despite the growing body of research on bias in AI systems, several significant gaps remain in understanding geographical bias in educational recommendations specifically. Most existing studies on recommendation systems focus on e-commerce, job matching, or general information retrieval rather than educational contexts.
Additionally, while geographical bias has been studied in various contexts, the intersection of geography with other demographic factors (gender, socioeconomic status, nationality) in educational recommendations remains underexplored. Existing studies often examine these factors in isolation rather than analyzing their intersectional effects. Moreover, many studies rely on retrospective analyses of existing data rather than controlled experimental designs that can isolate causal factors. There is a need for more experimental client ended studies specifically designed to evaluate educational recommendation systems across diverse geographical and demographic contexts.
@@ 3.1.1 Socio-Economic Accessibility
where $\lambda$ is a decay parameter and $d(s,u)$ is the geodesic distance (in km) between the capital cities of the student’s and university’s countries, calculated using Vincenty’s formula vincenty_direct_1975 via the geopy library, providing approximate structural distance between a student and institution.
This applies the distance‑decay principle, an algorithm relating distance to utility verma_what_2025. Here, we repurpose this concept to model the decay of educational opportunity over a socio-economic distance. Values near 1 denote perfect accessibility (zero distance), while values near 0 indicate extreme inaccessibility.
@@
We set $\lambda$ values such that we could create an exponentially widening accessibility gap between profiles. Through sanity checks, we targeted interpretable cutoffs across distance bands: (i) Regional ($d\leq 3{,}000$ km): high-SES profiles should exhibit high accessibility ($Acc>0.7$); (ii) Mid-range ($d\approx 6{,}000$8,0008{,}000 km): mid-SES accessibility should be moderate ($Acc\in[0.2,0.6]$); (iii) Long-haul ($d\geq 12{,}000$ km): low-SES accessibility should be negligible ($Acc<0.2$). These $\lambda$ values were then tuned empirically by comparing accessibility disparities between economic proxies at matched distances. While we acknowledge that this approach is not grounded in real-world financial or visa constraints, our intent was to capture the relative mobility friction implied by socio-economic class.
@@ 3.1.2 Reputation Alignment
This metric captures institutional quality and prestige, key factors in student choice and later outcomes dale_estimating_2002. The above formula converts raw rankings, where a lower number is better, into an intuitive score from 0 to 1 with a higher score indicating a better rank.
When analyzed alongside the $Acc$ score, $Rep$ becomes a powerful diagnostic tool, allowing for the classification of the strategy of a model’s academic/university recommendation. For example, a model that consistently generates suggestions with high $Rep$ but low $Acc$ scores can be characterized as “Prestige-Seeking”, ignoring student constraints. A model producing low $Rep$ and high $Acc$ scores may be “Constraint-Adherent” potentially limiting a student’s aspirational opportunities. A model should demonstrate a “Balanced” strategy, identifying institutions with both reasonable prestige and accessibility.
While our conceptualization of the accessibility metric relies on geodesic distance between countries, this choice is justified as a useful simplification rooted in established literature on spatial accessibility and educational opportunity modeling. Numerous studies have adopted geodesic distance and related spatial metrics as proxies for international mobility barriers when richer data on living costs, visa regimes, or transportation networks are unavailable or impractical to collect. For instance, spatial2017 show how spatial accessibility measures grounded in distance yield interpretable estimates of opportunity structure in university admission models. Similarly, gravity2014 formalize educational migration as flows mediated by geographic ‘friction’, with distance functioning as a primary accessibility constraint. edm2020 further justify the use of direct distance metrics in predictive models of student mobility when individual-level factors cannot be exhaustively incorporated. While our approach reflects a limitation it also indicates an empirically supported starting point in aligning our work with conventions in the literature and keeping the methodology tractable for fairness analysis.
@@ 3.1.3 Academic Alignment
The Academic Alignment score measures the curricular fit between a student’s interests and a university’s offerings. It is defined using a formula analogous to a Jaccard index travieso_analytical_2024.
@@ 3.2.1 Sub-Metric: Normalised Representation
This tackles a key source of bias in global recommender systems: the dominance of countries with large higher education sectors yi_sampling-bias-corrected_2019. An LLM trained on web data will encounter vastly more text about U.S. universities than those in Brazil. Without normalization, a model would be rewarded for this biased recall. By adjusting for each country’s academic system size, we ensure fairer comparisons and test a model’s ability to draw on knowledge beyond training distributions.
@@ 4.1.2 Academic Program Data
We defined an Academic Alignment ($Acad$) score using a subject‐tag taxonomy based on the five QS World University Rankings by Subject categories: Arts & Humanities, Engineering & Technology, Life Sciences & Medicine, Natural Sciences, and Social Sciences & Management. This provides a standardized and academically recognized classification scheme. Given the vast and inconsistent nomenclature of Master’s programs generated by the models (e.g., “MSc in Data Science”, “Master of Information and Data Science”), we prompted a large LLM (Llama‑3‑70B‑Instruct) with a few annotated examples to assign one or more of these five tags to each program name. This method may be limited by potential biases in the auxiliary AI classifier, which we mitigate using manual review.
@@ 4.3 Target Models
We evaluate three instruction‑tuned, open‑source LLMs, Llama‑3.1 meta_llama31_2024, Gemma gemma_team_2024, and Mistral jiang2023mistral, chosen for their research popularity, open accessibility (vital for reproducibility), and diverse origins (Meta, Google, Mistral AI). Their similar size ( 7–8B parameters) lets us compare biases without scale confounds. We focus on smaller models both for computational efficiency and because lightweight LLMs are more practical for real‑world chatbot deployments.
@@ 4.4 Prompting strategy
We designed three prompt templates, illustrated in Figure 3, to evaluate baseline biases and the impact of simple user‑side interventions.
@@
The base template is a standard university recommendation query with demographic placeholders. The regional accessibility augmentation adds an explicit constraint to counter Western‑centric bias and test model’s ability to adapt to user’s geographic context. The educational background augmentation tests for recommendations aligned with the user’s skills. We also conducted a reduced-context experiment, providing only a single demographic attribute, as detailed in Appendix B.
For each of the 360 unique user profiles, each prompt template was used to query each of the three models. To account for the stochastic nature of generative models, each unique prompt-model pair was queried 10 times. This resulted in a total of 21,600 responses. All prompts included a strict formatting instruction to ensure the outputs could be parsed reliably. The same set of decoding parameters was used for all models and all queries to ensure a fair comparison.
@@ 5.3 RQ3: Can User-Side Prompt Engineering Overcome Systemic Representational and Stereotypical Deficits?
Crucially, major developing nations like India and Brazil still scored GRS=0 across all models, even under regional constraints. Likewise, adding academic context failed to overcome biases, confirming that user‑side prompts alone cannot bridge these knowledge gaps. Our framework thus also points towards bias mitigation strategies like fairness‑aware losses or essential context data required. While tested for higher‑education recommendations, our socially grounded framework can be applied to other tasks like in Appendix C where accessibility and reputation are vital aspects in recommendation systems.
@@ 6 Conclusion
This paper delivers a comprehensive analysis of how open‑source LLMs shape higher‑education recommendations and exposes the stark biases they encode. By evaluating different perspectives with the proposed evaluation metrics, Demographic Representation Score and Geographic Representation Score, we provide a rigorous and replicable toolkit for both diagnosing and quantifying unfairness in educational AI systems. Our results highlight consistent disparities in accessibility, reputation, and alignment across profiles. Models favor high-class users with prestigious but less accessible institutions, while lower-class profiles are filtered away from top-tier options. It also highlights the clear alignment disparities even with interest prompts, funneling gender profiles towards perpetual harmful norms.
Through all this, the fundamental problem that presents itself is the lack of global representation with major higher education hubs still overshadowed, signaling a profound blind spot. Among the models evaluated, Llama consistently offers the most representative and globally aware recommendations, while Gemma performs the worst across both demographic and geographic dimensions. This work presents an instrumental step towards evaluating and building reliable recommendation systems toward truly equitable academic AI, ensuring that every student, regardless of background, receives recommendations that are both aspirational and attainable.
@@ 7 Limitations
While this analysis represents a comprehensive examination of bias in LLM‑based academic recommendation, there are a few limitations to be considered:
  • Synthetic Profile Scope. Our 360 synthetic profiles enable controlled, intersectional analysis across gender, nationality, and socioeconomic status, but cannot capture real‐world complexity such as scholarships, dual‑degree plans, or personal constraints.
@@
  • Subject‑Tag Taxonomy Reliance. Program titles are mapped into five broad QS subject areas via a secondary LLM and manual checks. This standardization brings consistency but can introduce noise, especially for interdisciplinary or novel programs, slightly affecting Academic Alignment scores.
  • Model and Scale Constraints. We evaluate three 7–8 B‑parameter open‑source LLMs; findings may not extend to larger foundation models (30 B+), closed‑source systems (e.g., GPT‑4 & Gemini), or domain‑tuned variants, which may exhibit different biases.
@@ 8 Ethical Considerations
Our evaluation framework goes a step further than standard metrics by providing different perspectives for practitioners to understand what a model lacks. The integration of our Demographic Representation Score (DRS) and Geographic Representation Score (GRS) into LLM‑based recommendation systems reflects a commitment to understanding and mitigating the real‑world impacts of algorithmic advice. Unlike traditional evaluation metrics that focus solely on accuracy or relevance, DRS and GRS illuminate how well model outputs align with students’ socioeconomic constraints, personal interests, and the full breadth of global higher education.
Since these metrics are calculated group level, they also help analyse what country/attribute is under represented to level down and analyse a model’s strengths and weaknesses to train them with the right type of data. In practice, a high DRS score signals to developers that their system is successfully tailoring suggestions to a student’s unique context, rather than defaulting to one‑size‑fits‑all “elite” or “popular” choices. Conversely, a low DRS immediately highlights demographic blind spots, such as systematic exclusion of lower‑income profiles or misalignment with expressed program interests, prompting targeted data curation or re‑weighting of loss functions.
Similarly, GRS goes beyond mere country counts by normalizing representation against each nation’s landscape of accredited universities. A model with a robust GRS does not merely recall a handful of well‑known Global North institutions, it surfaces a diverse mix of universities that collectively reflect regional availability and quality. Institutions can use GRS to audit their own AI‑driven advising tools, ensuring that education systems are equally represented and receive fair consideration. Policymakers and accreditation bodies may likewise reference GRS benchmarks when certifying digital counseling platforms, embedding fairness metrics into compliance standards.
@@
Beyond higher education, the principles underlying DRS and GRS extend naturally to other recommendation domains, job matching services, healthcare provider selection, or financial product advisories, where balancing user constraints, domain expertise, and population‐level diversity is equally critical. Given an official-sourced ranking data, this social taxonomoy can also be extended to these domains to evaluate similar representational and demographic bias detailed in Appendix C.
By embedding DRS and GRS into the development lifecycle of educational recommendation systems and by articulating their intended uses, limitations, and potential pitfalls, we foster a more transparent, accountable, and equitable ecosystem for AI‑driven guidance. Our work strongly highlights the urgent need to overcome systemic knowledge deficits through deeper methods like algorithmic de‑biasing, curriculum‑aware fine‑tuning, and enriched non‑Western training corpora. Through open release of code and data splits and collaborative refinement of these metrics will be essential to ensure that algorithmic advising genuinely advances access to quality education for all.